OpenStudy (anonymous):

Find the area between the curves: y=√x and y=x^2 Please show steps!

OpenStudy (anonymous):

First find the intersection of the two functions: \[\sqrt{x}=x^{2}\] x=0 and 1 then integrate the difference of the two functions noting the one on top and the one on the bottom: \[\int\limits_{0}^{1}\sqrt{x}-x^{2}dx\] this works out to be:1/3